2(x + 3) – 4 = 12 + (x - 3)
Answer by HayleyAlyce(8)

2(x + 3) – 4 = 12 + (x - 3)
2x+6-4=12+x-3
2x+2=9+x
-2 -2
2x=7+x
-x -x
x=7